solve
A) -7a+17=-18
B) 3/4=x/2+5/8
C) x-5>-4
D) 3x/-7<-3

Answer by checkley77(12844)

A) -7a+17=-18
-7a=-18-17
-7a=-35
a=-35/-7
a=5 ans.
B) 3/4=x/2+5/8
x/2=3/4-5/8
x/2=(2*3-5)/8
x/2=(6-5)/8
x/2=1/8
8x=2
x=2/8
x=1/4 ans.
C) x-5>-4
x>-4+5
x>1 ans.
D) 3x/-7<-3
3x>-7*-3 [note when multiplying by a negative value the sign.]
3x>21
x>21/3
x>7 ans.